# Integrating Spatiotemporal Features in LSTM for Spatially Informed COVID-19 Hospitalization Forecasting

### v1

@@ Abstract
The COVID-19 pandemic’s severe impact highlighted the need for accurate, timely hospitalization forecasting to support effective healthcare planning. However, most forecasting models struggled, especially during variant surges, when they were needed most. This study introduces a novel Long Short-Term Memory (LSTM) framework for forecasting daily state-level incident hospitalizations in the United States. We present a spatiotemporal feature, Social Proximity to Hospitalizations (SPH), derived from Facebook’s Social Connectedness Index to improve forecasts. SPH serves as a proxy for interstate population interaction, capturing transmission dynamics across space and time. Our parallel LSTM architecture captures both short- and long-term temporal dependencies, and our multi-horizon ensembling strategy balances consistency and forecasting error. Evaluation against COVID-19 Forecast Hub ensemble models during the Delta and Omicron surges reveals superiority of our model. On average, our model surpasses the ensemble by 27, 42, 54, and 69 hospitalizations per state on the $7^{th}$, $14^{th}$, $21^{st}$, and $28^{th}$ forecast days, respectively, during the Omicron surge. Data-ablation experiments confirm SPH’s predictive power, highlighting its effectiveness in enhancing forecasting models. This research not only advances hospitalization forecasting but also underscores the significance of spatiotemporal features, such as SPH, in refining predictive performance in modeling the complex dynamics of infectious disease spread.
@@
  • Deep learning can be used to reliably forecast the spread of infectious diseases.
@@ 1 Introduction
Since the first laboratory-confirmed case of COVID-19, an infectious disease caused by the SARS-CoV-2 virus, in the United States in January 2020, there have been approximately 111 million total confirmed infections, 1.2 million deaths, and 7 million hospitalizations as of June 2023. Despite the end of the pandemic phase and declining mortality rates, COVID-19 remains a significant global health concern. According to the CDC COVID-19 Dashboard, as of 2024, the disease exhibited a peak weekly test positivity of 18% in the US. Furthermore, while the recorded 4.8 hospitalization rate per 10,000 population as of August 10, 2024, may appear comparatively low, it underscores the continuing impact of the disease. According to communications received from the CDC, hospitals will be mandated to report COVID-19 hospitalizations again starting in mid-November 2024, indicating the resurgence of the disease. The COVID-19 pandemic strained healthcare resources and overloaded hospitals, exacerbating the dramatic loss of human life. SARS-CoV-2 spreads rapidly, causing severe complications due to its high reproduction rate, ability to spread via asymptomatic individuals, prevalence in close-contact settings in densely populated areas, mutation into more transmissible variants, and inconsistent application of preventive public health measures across the US. As a result, the demand for travel nurses surged during the pandemic, aligning with shifting of COVID-19 infection hotspots (Cole et al. 2021, Longyear et al. 2020). This was partially a geospatial problem related to the timely allocation of limited human and medical resources. Reliable geographic forecasting of COVID-19 hospital admissions could have alleviated this burden through policy-relevant decision-making and proactive allocation of resources in regional hotspots, i.e.  areas of increased demand.
@@
Accurately forecasting the geographic spread of COVID-19, particularly hospitalizations, presented significant challenges. Early in the pandemic, the time between symptom onset and hospitalization varied between 3 and 10.4 days, averaging about 5 days (Faes et al. 2020, Galmiche et al. 2023), which provided advance albeit short notice for hospitals to prepare for an upcoming influx of inpatients after observing a rise in test positivity rates. However, with the emergence of variants, such as Omicron (B.1.1.529), the time lag between the surge in cases and hospitalizations shortened to around 3.6 days on average (Jansen 2021, Song et al. 2022, Grant et al. 2020, Tanaka et al. 2022, Galmiche et al. 2023), reducing the preparation window for healthcare providers and complicating the forecasting efforts. Additionally, the rise in at-home testing contributed to a significant underreporting of cases (Liu et al. 2021, Irons and Raftery 2021), rendering case reports unreliable and underscoring the importance of directly forecasting hospitalizations, which are mandated to be reported by the Centers for Medicare and Medicaid Services (CMS).
@@
Lastly, unlike the previous surges, hospitalization incidence largely decoupled from case incidence during the Omicron surge (Moore and Baden 2022). Although the Omicron variant infected a far larger share of the population more rapidly than previous variants, it did not result in a similar proportion of hospitalizations. This could be due to the immunity granted by vaccination, prior infections, and potentially the characteristics of the Omicron variant itself. Therefore, direct forecasting of hospitalizations, rather than as a proportion of confirmed cases, is critical for planning interventions and resource allocation. Additionally, a model will need to balance learning from short- and long-term changes in hospitalization rates to perform optimally in the face of new variants. Moreover, to improve hospitalization forecasting, the disease spread dynamics should be taken into account. SARS-CoV-2 is an airborne virus, and therefore, human interaction leads to disease diffusion in space and time, and results in spatiotemporal hotspots. While infections in a county may be influenced by social connections, some of those result in hospitalizations, leading to additional regional spillover effects. This means infection and hospitalization in one spatial jurisdiction can affect the trends in connected jurisdiction (Wang and Aydin 2020) and successful incorporation of spatial dependencies and connectivity is a key factor in improving the quality of forecasts (Lucas et al. 2023).
@@
To address these challenges, we develop a novel hospitalization forecasting model based on deep learning with spatiotemporal features derived from social media friendship data. We leverage Facebook’s Social Connectedness Index (SCI), which is derived from aggregate social media friendship counts across spatial units, to capture potential interstate transmission. Our model operates at the spatial unit level, and not at the individual level, which mitigates concerns related to individual usage of Facebook. The high correlation of regional COVID-19 caseloads with SCI values has been shown in previous studies (Kuchler et al. 2022), and another study showed that SCI is more predictive than SafeGraph’s cell-phone-based mobility data, albeit with a simpler tree-based ensemble model (Vahedi et al. 2021). While we acknowledge that no dataset is fully representative, Facebook’s SCI data remains one of the most comprehensive resources available for measuring regional connectivity and their impact on disease spread.
@@
    Development of novel interstate spatiotemporal features using social media-derived aggregate data to incorporate spatial dependence and improve performance and reduce model variance.
@@
    Development of a novel multi-horizon ensembling strategy to balance between predictive performance and consistency in output sequences.
@@
    Integration of quantile loss for probabilistic forecasts for each state to characterize uncertainty and evaluation using the weighted interval score, which jointly penalizes dispersion, under- and over-prediction.
We rigorously evaluate the developed spatiotemporal feature, multi-horizon model performance, uncertainty, and compare against baseline models on data collected during the Omicron surge in the U.S., as the most challenging period for hospitalization forecasting during the pandemic. We also release open-source code, models, and data processing pipelines for reproducibility and reuse by the research community, and for application to other infectious diseases.
@@ 2 Background
A comprehensive review of COVID-19 modeling identified approximately 22,000 publications related to the topic, spanning various modeling approaches, notably compartmental and statistical models(Cao and Liu 2022).
The United States COVID-19 Forecast Hub (hereafter referred to as the Hub) is an online repository created by the US Centers for Disease Control and Prevention (CDC) and an academic research lab at the University of Massachusetts Amherst. Launched in April 2020 (Cramer et al. 2022a), the repository fulfills several crucial roles in monitoring COVID-19 in the US, including providing reported data of incident cases and deaths from Johns Hopkins University (JHU)’s Center for System Science and Engineering (CSSE) Dataset (Dong et al. 2020) and incident hospitalizations from HealthData.gov [^1] https://healthdata.gov/Hospital/COVID-19-Reported-Patient-Impact-and-Hospital-Capa/g62h-syeh. The COVID-19 dataset collected during the pandemic is uniquely large with high spatial coverage, and well quality controlled, thanks to the stringent legal and regulatory mandates. This has provided an excellent benchmark dataset for evaluating the performance of new modeling approaches, one that we leverage here. Having a central point of data acquisition has proven a valuable resource for researchers. Additionally, the repository gathers and publishes point and probabilistic forecasts of incident cases, deaths, hospitalizations, and accumulative deaths due to COVID-19 at the national, state, and county levels in the U.S. submitted by different research groups employing various methodologies. These forecasts facilitate comparative analyses of model performance and the development of ensemble models by combining forecasts from multiple models (see section 2.1).
Hospitalization forecasts submitted to the Hub vary in time horizon and resolution, ranging from 1 to 130 days ahead, with longer horizons being the focus of scenario modeling. In this study, we primarily focus on short-term (28-day) incident hospitalization forecasts, the Hub-standard focus for hospitalization forecasting. Although county-level forecasting can provide finer-grained insights, the CDC and the Hub standardized state-level hospitalization forecasts as their operational benchmark (Cramer et al. 2022a). This decision was partially driven by the fact that county-level hospitalization data were not consistently available nationwide during the evaluation period. Moreover, focusing on state-level data ensures consistency and comparability across different research teams submitting forecasts to the Hub, while aligning with practical constraints of reporting and data reliability. Beyond point forecasts, the Hub mandates the submission of quantiles from 23 probability levels (including 0.01, 0.025, 0.05, 0.10, 0.15, 0.20, 0.25, 0.30, 0.35, 0.40, 0.45, 0.50, 0.55, 0.60, 0.70, 0.75, 0.80, 0.90, 0.95, 0.975, 0.99) for state-level incident hospitalization forecasts.
@@ 2.1 COVID Forecast Hub Ensemble
The COVID Forecast Hub uses the submitted forecasts by contributing teams, and generates three distinct models: a baseline model for reference and two ensemble models, all of which we incorporate for comparison in our experiments. The COVIDhub-baseline model operates as a persistence model for reference (Cramer et al. 2022a), where the point forecast for any future time is equal to the most recently observed value. Ensemble methods, a meta-learning approach, combines predictions from multiple models to enhance predictive accuracy. Starting from the week of November 15, 2021, COVIDhub-4_week_ensemble computes an equally weighted median of eligible forecasts for cases and hospitalizations. While the ensemble may not consistently yield the most accurate forecast for individual dates, it demonstrates stable forecasting performance over time compared to other models within the Hub (Bracher et al. 2021). In contrast, the forecasts generated by the COVIDhub-trained_ensemble are a weighted median of forecasts from the top-performing individual models, ranked based on relative Weighted Interval Score (WIS) assessed on performance for previous forecasts submissions (Bracher et al. 2021).
@@ 2.2 Compartmental Models
Among the 17 models in the Hub that include incident hospitalization forecasting, seven utilize compartmental models, a traditional approach with a history of application in modeling infectious diseases. The general idea of compartmental models is dividing the population into compartments throughout the progression of the disease (Kendall 1956). The most basic model, SIR, divides people into Susceptible, Infected, and Recovered groups. The dynamics of an epidemic without considering the dynamics of birth and death can be expressed by the following ordinary differential equations:
@@
The more popular SEIR model extends the SIR model by introducing an Exposed compartment, reflecting an incubation period where individuals are exposed to the pathogen but before the onset of symptoms. This modification is encapsulated in the modified equations 1b and 1c, now referred to as Equations (2b) and (2c), respectively:
@@
A common criticism of the basic compartmental models is that they assume spatiotemporal homogeneity of the spatial units, and oversimplify the complex disease processes (Ansumali and Prakash 2020, Getz et al. 2019). The complex and dynamic patterns of COVID-19 transmission require updating the disease parameters with epidemiological expert supervision in compartmental models. SEIR models can be extended to better incorporate spatiotemporal heterogeneity and human mobility, enhancing their applicability to dynamic epidemic conditions. For instance, the JHUAPL-Bucky model integrates spatial mobility matrices and age-based contact matrices to capture interactions across regions and different age demographics (Cramer et al. 2022a). Similarly, models like GLEAM developed by Balcan et al. utilize mobility networks and spatial diffusion to model disease spread across geographic regions to account for local variations in population density and movement patterns. These models, including USC’s SI-kJalpha, which adjusts for heterogeneous infection rates due to variants and vaccination statuses (Srivastava et al. 2020), attempt to capture complex realities of infectious disease transmission within compartmental frameworks. Moreover, recent developments like the Multistrain SEIR Model by Laaroussi et al. and Seibel et al. integrate distributed vaccination policies and behavioral heterogeneity (which requires additional data availability), respectively, incorporating spatial and human factors influencing disease dynamics. The JHUAPL-Bucky, GLEAM, and SI-kJalpha are included in the COVIDhub-4_week_ensemble, and due to their high performance, they are also included in the COVIDhub-trained_ensemble, against which we compare our models in this paper. However, it is worth noting that deep learning-based models allow for more complex modeling through deeper networks with non-linear activation functions, with orders of magnitude more learnable parameters to capture heterogeneous dynamics.
@@ 2.3 Statistical Models
Due to the availability and high spatiotemporal resolution of COVID-19 incidence reports, statistical approaches have grown in popularity and demonstrated their performance in accurate forecasting during the COVID-19 pandemic(Reinhart et al. 2021, Rahimi et al. 2023, Vaughan et al. 2023). Where compartmental models make assumptions of population and epidemic characteristics, statistical approaches can learn complex patterns directly from observed data. However, these data-driven methods are highly dependent on data quality and usually have poor interpretability and scenario-building capability.
As COVID-19 cases and hospitalizations are highly temporally correlated, many teams chose time-series forecasting methods among the big family of statistical approaches. The Auto-Regressive Integrated Moving Average (ARIMA) model is a time series forecasting method that makes future predictions by analyzing historical data. Some teams applied such methods to forecast incident cases and hospitalizations and reach satisfactory prediction accuracy (Benvenuto et al. 2020, Alabdulrazzaq et al. 2021). Instead of using a single autoregressive method, CMU Delphi Group used a cumulative distribution function-space-averaged ensemble of three models, including a simple autoregressive model for point prediction, a quantile autoregressive model with additional case covariates, and a direction-stratified quantile autoregressive model (Reinhart et al. 2021).
Deep learning-based approaches have shown great potential in epidemiological modeling in recent years. GT-DeepCOVID model is a deep learning-based approach that learns the dependence of hospitalization and mortality rate from syndromic, demographic, mobility, and clinical data (Rodriguez et al. 2021). To address the poor interpretability, the model also integrates an explainability module to discover the impact of different variables. Long Short-Term Memory (LSTM) is an alternate deep learning-based approach that is more skillful at capturing long-distance dependencies in time series data. LSTM and its variants were applied in COVID-19 forecasting with COVID-19 time series data as input (Lucas et al. 2023).

### v2

@@ Abstract
The COVID-19 pandemic’s severe impact highlighted the need for accurate and timely hospitalization forecasting to support effective healthcare planning. However, most forecasting models struggled, particularly during variant surges, when they were most needed. This study introduces a novel parallel-stream Long Short-Term Memory (LSTM) framework to forecast daily state-level incident hospitalizations in the United States. Our framework incorporates a spatiotemporal feature, Social Proximity to Hospitalizations (SPH), derived from Meta’s Social Connectedness Index, to improve forecasts. SPH serves as a proxy for interstate population interaction, capturing transmission dynamics across space and time. Our architecture captures both short- and long-term temporal dependencies, and a multi-horizon ensembling strategy balances forecasting consistency and error. An evaluation against the COVID-19 Forecast Hub ensemble models during the Delta and Omicron surges reveals the superiority of our model. On average, our model surpasses the ensemble by 27, 42, 54, and 69 hospitalizations per state at the 7-, 14-, 21-, and 28-day horizons, respectively, during the Omicron surge. Data-ablation experiments confirm SPH’s predictive power, highlighting its effectiveness in enhancing forecasting models. This research not only advances hospitalization forecasting but also underscores the significance of spatiotemporal features, such as SPH, in modeling the complex dynamics of infectious disease spread.
@@
  • Deep learning can be used to more reliably forecast the spread of infectious diseases.
@@ 1 Introduction
Since the first laboratory-confirmed case of COVID-19, an infectious disease caused by the SARS-CoV-2 virus, in the United States (U.S.) in January 2020, there have been approximately 111 million total confirmed infections, 1.2 million deaths, and 7 million hospitalizations as of June 2023. Despite the end of the pandemic phase and declining mortality rates, COVID-19 remains a significant global health concern. According to the Centers for Disease Control and Prevention (CDC) COVID-19 Dashboard, the disease exhibited a peak weekly test positivity of 18% in the U.S. in 2024. Although the recorded hospitalization rate of 4.8 per 10,000 population on August 10, 2024, may appear comparatively low, it underscores the continuing impact of the disease. According to communications received from the CDC, hospitals are mandated to report COVID-19 hospitalizations again starting in mid-November 2024, indicating the resurgence of the disease. The COVID-19 pandemic strained healthcare resources and overloaded hospitals, exacerbating the dramatic loss of human life. SARS-CoV-2 spreads rapidly, causing severe complications due to its high reproduction rate, the ability to spread via asymptomatic individuals, the prevalence of close-contact settings in densely populated areas, continual mutation into more transmissible variants, and the inconsistent application of preventive public health measures across the U.S. As a result, the demand for travel nurses surged during the pandemic, aligning with shifts in COVID-19 infection hotspots (Cole et al. 2021, Longyear et al. 2020). This was partially a geospatial problem related to the timely allocation of limited human and medical resources. Reliable geographic forecasting of COVID-19 hospital admissions could have alleviated this burden through policy-relevant decision-making and proactive allocation of resources in regional hotspots (i.e.  areas of increased demand).
@@
Accurately forecasting the geographic spread of COVID-19, particularly hospitalizations, presented significant challenges. Early in the pandemic, the time between symptom onset and hospitalization varied from 3 to 10.4 days, averaging about 5 days (Faes et al. 2020, Galmiche et al. 2023); this interval provided advance, albeit short, notice for hospitals to prepare for an upcoming influx of inpatients after observing a rise in test positivity. However, with the emergence of variants, such as Omicron (B.1.1.529), the lag between the surge in cases and hospitalizations shortened to around 3.6 days on average (Jansen 2021, Song et al. 2022, Grant et al. 2020, Tanaka et al. 2022, Galmiche et al. 2023), reducing the preparation window for healthcare providers and complicating forecasting efforts. Additionally, the rise of at-home testing contributed to significant underreporting of cases (Liu et al. 2021, Irons and Raftery 2021), rendering case reports unreliable and reinforcing the importance of forecasting hospitalizations directly, which are mandated to be reported by the Centers for Medicare and Medicaid Services (CMS).
@@
Lastly, unlike previous surges, hospitalization incidence largely decoupled from case incidence during the Omicron surge (Moore and Baden 2022). Although the Omicron variant infected a much larger share of the population more rapidly than earlier variants, it did not result in a similar proportion of hospitalizations. This could be due to the immunity granted by vaccination, prior infections, and potentially the characteristics of the Omicron variant itself. Therefore, direct forecasting of hospitalizations, rather than a proportion of confirmed cases, is critical for intervention planning and resource allocation. Additionally, a model must balance learning from short- and long-term changes in hospitalization rates to perform optimally in the face of new variants. Moreover, to improve hospitalization forecasting, the disease spread dynamics should be taken into account. SARS-CoV-2 is an airborne virus; therefore, human interaction leads to disease diffusion in space and time, resulting in spatiotemporal hotspots. While infections in a county may be influenced by social connections, some of those result in hospitalizations, leading to additional regional spillover effects. This means that infection and hospitalization in one spatial jurisdiction can affect trends in connected jurisdictions (Wang and Aydin 2020), and successful incorporation of spatial dependencies and connectivity is a key factor in improving the quality of forecasts (Lucas et al. 2023).
@@
To address these challenges, we developed a novel hospitalization forecasting model based on deep learning with spatiotemporal features derived from social media friendship data. Our model leverages Meta’s Social Connectedness Index (SCI), which is derived from aggregated social media friendship counts across spatial units, to capture potential interstate transmission. Because our model operates at the spatial unit level, rather than the individual level, it mitigates concerns related to individual Facebook usage. The high correlation of regional COVID-19 caseloads with SCI values has been shown in previous studies (Kuchler et al. 2022), and another study found that SCI is more predictive than SafeGraph’s cell-phone-based mobility data, albeit with a simpler tree-based ensemble model (Vahedi et al. 2021). Although we acknowledge that no dataset is fully representative, Facebook’s SCI data remains one of the most comprehensive resources available for measuring regional connectivity and its impact on disease spread.
@@
    Development of novel interstate spatiotemporal features using social media-derived aggregate data to incorporate spatial dependence, improve performance, and reduce model variance.
@@
    Design of a multi-horizon ensembling strategy that balances between predictive performance and consistency in output sequences.
@@
    Integration of quantile loss for probabilistic forecasts for each state to characterize uncertainty and evaluation with the weighted interval score, which jointly penalizes dispersion, under- and over-prediction.
This paper also reports rigorous evaluations of our proposed spatiotemporal features, multi-horizon model performance, uncertainty, and compares our model against baseline models on data collected during the Omicron surge in the U.S.—the most challenging period for hospitalization forecasting during the pandemic, as well as the earlier Delta surge period. We also release open-source code, models, and data processing pipelines for reproducibility and reuse by the research community, as well as for application to other infectious diseases.
@@ 2 Background
A comprehensive review of COVID-19 modeling identified approximately 22,000 publications related to the topic, spanning various modeling approaches, most notably compartmental and statistical models (Cao and Liu 2022).
The United States COVID-19 Forecast Hub (hereafter referred to as the Hub) is an online repository created by the U.S. Centers for Disease Control and Prevention (CDC) and an academic research lab at the University of Massachusetts Amherst. Launched in April 2020 (Cramer et al. 2022a), the repository fulfills several crucial roles in monitoring COVID-19 in the U.S., including providing reported data on incident cases and deaths from Johns Hopkins University (JHU) Center for System Science and Engineering (CSSE) Dataset (Dong et al. 2020) and incident hospitalizations from HealthData.gov [^1] https://healthdata.gov/Hospital/COVID-19-Reported-Patient-Impact-and-Hospital-Capa/g62h-syeh. The COVID-19 dataset collected during the pandemic is uniquely large with high spatial coverage, and well quality controlled, thanks to the stringent legal and regulatory mandates. This has provided an excellent benchmark dataset for evaluating the performance of new modeling approaches, one that we leverage here. Having a central point of data acquisition has proven to be a valuable resource for researchers. Additionally, the repository collects and publishes point and probabilistic forecasts of incident cases, deaths, hospitalizations, and cumulative deaths due to COVID-19 at national, state, and county levels in the U.S. submitted by various research groups employing different methodologies. These forecasts facilitate comparative analyses of model performance and the development of ensemble models by combining forecasts from multiple models (see section 2.1).
Hospitalization forecasts submitted to the Hub vary in time horizon and resolution, ranging from 1 to 130 days ahead, with longer horizons being the focus of scenario modeling. In this study, we primarily focus on short-term (28-day) incident hospitalization forecasts—the Hub-standard focus for hospitalization forecasting. Although county-level forecasting can provide finer-grained insights, the CDC and the Hub standardized state-level hospitalization forecasts as their operational benchmark (Cramer et al. 2022a). This decision was partly driven by the fact that county-level hospitalization data were not consistently available nationwide during the evaluation period. Moreover, focusing on state-level data ensures consistency and comparability across different research teams submitting forecasts to the Hub, while aligning with practical constraints of reporting and data reliability. Beyond point forecasts, the Hub mandates the submission of quantiles from 23 probability levels (including 0.01, 0.025, 0.05, 0.10, 0.15, 0.20, 0.25, 0.30, 0.35, 0.40, 0.45, 0.50, 0.55, 0.60, 0.70, 0.75, 0.80, 0.90, 0.95, 0.975, 0.99) for state-level incident hospitalization forecasts.
@@ 2.1 COVID Forecast Hub Ensemble
The Hub uses the submitted forecasts from contributing teams to generate three distinct models: a baseline model for reference and two ensemble models, all of which are incorporated for comparison in our experiments. The COVIDhub-baseline model operates as a persistence model for reference (Cramer et al. 2022a), where the point forecast for any future time is equal to the most recently observed value. Ensemble methods, a meta-learning approach, combine predictions from multiple models to enhance predictive accuracy. Starting from the week of November 15, 2021, COVIDhub-4_week_ensemble computes an equally weighted median of eligible forecasts for cases and hospitalizations. Although the ensemble may not consistently yield the most accurate forecast for individual dates, it demonstrates stable forecasting performance over time compared to other models within the Hub (Bracher et al. 2021). By contrast, the forecasts generated by the COVIDhub-trained_ensemble are a weighted median of forecasts from the top-performing individual models, ranked based on the relative Weighted Interval Score (WIS) assessed on performance for previous forecast submissions (Bracher et al. 2021).
@@ 2.2 Compartmental Models
Among the 17 models in the Hub that forecast incident hospitalizations, seven employ compartmental models, a traditional approach with a history of application in modeling infectious diseases. The general idea of compartmental models is to divide the population into compartments throughout the progression of the disease (Kendall 1956). The most basic model, SIR, divides people into Susceptible, Infected, and Recovered groups. The dynamics of an epidemic without considering the dynamics of birth and death can be expressed by the following ordinary differential equations:
@@
The more popular SEIR model extends the SIR model by introducing an Exposed compartment, which reflects the incubation period during which individuals are exposed to the pathogen but have not yet developed symptoms. This modification is encapsulated in the modified equations 1b and 1c, now referred to as Equations (2b) and (2c), respectively:
@@
A common criticism of basic compartmental models is that they assume spatiotemporal homogeneity of the spatial units, and oversimplify the complex disease processes (Ansumali and Prakash 2020, Getz et al. 2019). The complex and dynamic patterns of COVID-19 transmission require updating the disease parameters with epidemiological expert supervision in compartmental models. SEIR models can be extended to better incorporate spatiotemporal heterogeneity and human mobility, enhancing their applicability to dynamic epidemic conditions. For instance, the JHUAPL-Bucky model integrates spatial mobility matrices and age-based contact matrices to capture interactions across regions and different age demographics (Cramer et al. 2022a). Similarly, models like GLEAM developed by Balcan et al. utilize mobility networks and spatial diffusion to model disease spread across geographic regions to account for local variations in population density and movement patterns. These models, including USC’s SI-kJalpha, which adjusts for heterogeneous infection rates due to variants and vaccination statuses (Srivastava et al. 2020), attempt to capture complex realities of infectious disease transmission within compartmental frameworks. Moreover, recent developments, such as the Multistrain SEIR Model by Laaroussi et al. and Seibel et al., integrate distributed vaccination policies and behavioral heterogeneity (which require additional data availability), respectively, incorporating spatial and human factors that influence disease dynamics. The JHUAPL-Bucky, GLEAM, and SI-kJalpha are included in the COVIDhub-4_week_ensemble, and due to their high performance, they are also included in the COVIDhub-trained_ensemble, against which we compare our models in this paper. However, it is worth noting that deep learning-based models allow for more complex modeling through deeper networks with non-linear activation functions, with orders of magnitude more learnable parameters to capture heterogeneous dynamics.
@@ 2.3 Statistical Models
Due to the availability and high spatiotemporal resolution of COVID-19 incidence reports, statistical approaches have grown in popularity and demonstrated their performance in accurate forecasting during the COVID-19 pandemic (Reinhart et al. 2021, Rahimi et al. 2023, Vaughan et al. 2023). Whereas compartmental models make assumptions about population and epidemic characteristics, statistical approaches can learn complex patterns directly from observed data. However, these data-driven methods are highly dependent on data quality and usually have poor interpretability and scenario-building capability.
As COVID-19 cases and hospitalizations are highly temporally correlated, many teams have chosen time-series forecasting methods within the family of statistical approaches. The Auto-Regressive Integrated Moving Average (ARIMA) model is a time-series forecasting method that makes future predictions by analyzing historical data. Some teams have applied such methods to forecast incident cases and hospitalizations, achieving satisfactory prediction accuracy (Benvenuto et al. 2020, Alabdulrazzaq et al. 2021). Instead of using a single autoregressive method, CMU Delphi Group used a cumulative distribution function-space-averaged ensemble of three models, including a simple autoregressive model for point prediction, a quantile autoregressive model with additional case covariates, and a direction-stratified quantile autoregressive model (Reinhart et al. 2021).
Deep learning-based approaches have shown great potential in epidemiological modeling in recent years. The GT-DeepCOVID model is a deep learning-based approach that learns the dependence of hospitalization and mortality rates from syndromic, demographic, mobility, and clinical data (Rodriguez et al. 2021). To address its poor interpretability, the model also integrates an explainability module to discover the impact of different variables. Long Short-Term Memory (LSTM) is an alternative deep learning-based approach that is more skillful at capturing long-distance dependencies in time-series data. LSTM and its variants have been applied in COVID-19 forecasting with COVID-19 time series data as input (Lucas et al. 2023).